# Towards a turnkey approach to unbiased Monte Carlo estimation of smooth functions of expectations

### v2

@@ 4.1. Toy latent variable model
We set $\theta=1$ and sample a data set of size $n=10^{3}$. The goal is to unbiasedly estimate the log-likelihood $\sum_{i=1}^{n}\log p(y_{i}|\theta)$ for $\theta=1$ (a quantity which we are able to compute exactly for this model).
@@
$I_{j}:=\log\frac{1}{2^{j}}\sum_{h=1}^{2^{j}}\frac{p(Z_{h},y|\theta)}{q(Z_{h};y)}=\log\frac{1}{2^{j}}\sum_{h=1}^{2^{j}}w(Z_{h};y,\theta),\qquad Z_{1},\dots,Z_{2^{j}}\sim q(\cdot;y).$
and $\Delta_{k}:=I_{j+k+1}-(I_{j+k}^{E}+I_{j+k}^{O})/2$; $I_{j+k}^{E}$ (resp. $I_{j+k}^{O}$) being the IWAE estimate that only uses the even-indexed (resp. odd-indexed) samples. We consider the same proposal as in Shi and Cornish, (2021), i.e. $\mathcal{N}(\cdot;(y+\theta^{\star}\textsf{1}_{d})/2,\frac{2}{3}\textsf{Id}_{d})$, where $\theta^{\star}$ denotes the maximum likelihood estimator for $\theta$, which is shown to be optimal within a certain family in Rainforth et al., (2018).
The expected CPU cost (i.e. the number of generated $X_{j}$) is $\mathbb{E}[R]=1/p-1$ for our Taylor-based estimators, and $\mathbb{E}[2^{j+1+\widetilde{R}}]$ for the MLMC approach of Shi and Cornish, (2021). We adjust $p$ and $j$ so that both expectations are equal to a certain pre-specified budget $C$. We consider two values for $C$, namely $C=6$ and $C=96$.
@@
This experiment shows that the cost of MLMC has large empirical variance. In fact, the true variance of the cost is infinite (see Appendix E.1 for a proof of this statement). In contrast, the cost of the simple and cycling estimators remains close to its expectation.
The simple and cycling estimators outperform MLMC in terms of accuracy, with the simple estimator having larger variance than the cycling estimator for higher computational cost. This is particularly evident in the first plot of the second row, in which we compare the estimates for $\mathbb{E}[R]=96$, in this case the cloud of estimates obtaining with cycling is considerably more concentrated around the true value (the horizontal line). This is consistent with the theoretical results established in Section 2 which show that the variance of the simple estimator does not decay to 0 when the computational cost increases.
@@
As a note of caution, our Taylor-based estimators may give very poor results for a larger dimension $d$; see Appendix E.1 for extra results when $d=20$. Simple calculations show that, in that case, one would need to take $\mathbb{E}[R]\gg 10^{8}$ to ensure that $p<1-\beta^{2}$ (so that the estimators have finite variance). In this case, the inputs have very large variance, and presumably the only way to fix this issue is to come up with a different strategy to unbiasedly estimate $p(y|\theta)$, so that the corresponding estimates have much lower variance.
@@ 4.2. Independent component analysis
To estimate $\theta$, Allassonnière and Younes, (2012) use a stochastic approximation expectation maximisation (SAEM) algorithm. We will instead use stochastic gradient descent (Robbins and Monro,, 1951), implemented with our unbiased estimators of $\nabla_{\theta}f(m(\theta))=\nabla_{\theta}\log p(\theta|y)$. We call our two implementations simple-SGD and cycling-SGD. We also implement SGD with the gradient unbiasedly estimated using the MLMC strategy of Shi and Cornish, (2021) (MLMC-SGD). The bivariates $(X_{i},G_{i})$ (for $i=1,2,\ldots$) of Section 2.6 are generated using importance sampling; and the same importance sampling proposal is used for MLMC. For MLMC, we adjust its parameters $p$ and $j$ (defined in (17) with $\tilde{R}\sim\mathrm{Geometric}(p)$) so that $\mathbb{E}\left[2^{j+1+\widetilde{R}}\right]=n_{0}$. Similarly, $p$ for the Geometric random truncation of the simple and cycling gradient estimates is found by setting $\mathbb{E}\left[R\right]=n_{0}$. See Appendix E.2 for all the details.
@@
SAEM $7.20\cdot 10^{-7}$ 0.21 33
MLMC $3.38\cdot 10^{-6}$ 0.30 12 22400 9
simple $1.69\cdot 10^{-5}$ 0.28 10 90 15
cycling $1.63\cdot 10^{-4}$ 0.27 10 90 19
@@ 4.3. Exponential random graphs
We construct an unbiased estimates of $1/\mathcal{Z}(\theta)$, for any $\theta$, which may then be combined with the remaining tractable terms of (19) to perform Bayesian inference and model choice. We consider the same example as in Caimo and Friel, (2011): an $n=16$ node graph; the statistic is $s(y)=(\sum_{i<j}y_{ij},\sum_{i<j<k}y_{ij}y_{jk})$, which counts the number of edges and of two stars; $\theta\in{}^{2}$ is assigned the prior $N_{2}(0_{2},30I_{2})$. The data – the observed connectivity – is the Florentine family business graph (Padgett and Ansell,, 1993).
@@
4.3.1. Variance Comparison (fixed $\theta$)
We first fix $\theta=(-3.15,0.58)$ (a value in the high probability region of the posterior), and compare the performance of the simple and the cycling estimator.
We select the value of $x_{0}$ using the strategy in Section 3.1 with $n_{0}=10$. Once $x_{0}$ has been selected, we identify the corresponding $\widehat{\beta^{2}}$ as discussed in Section 3.2. This gives us a range for $p\in(0,1-\widehat{\beta^{2}})$. In the moderate variance regime we obtain $p\in(0,0.7)$ while for the low variance regime we have $p\in(0,0.95)$. In both cases, it is sufficient to set $\mathbb{E}[R]=1$ (i.e. $p=0.5$) to ensure that both the simple and the cycling estimator have finite variance.
Since the cost of obtaining the estimates of $\mathcal{Z}(\theta)$ is significantly larger than that of computing the estimates, selecting $\mathbb{E}[R]=1$ might be wasteful as our theoretical results in Section 2 guarantee that the cycling estimator has a lower variance than the simple estimator for large values of $\mathbb{E}[R]$. Thus, we also consider $\mathbb{E}[R]=10=n_{0}$ in our simulations. (Recall our recommendation in Section 3.2 is to set $\mathbb{E}[R]$ to $n_{0}$.)
Table 1 compares the simple and cycling estimators in these four different scenarios (two variance regimes, two values of $\mathbb{E}[R]$) according to two metrics (estimated on $10^{3}$ independent runs): work-normalised variance (relative variance times CPU cost), and proportion of negative $1/\mathcal{Z}(\theta)$ estimates. The results show that (a) the cycling estimator systematically outperforms the simple one; and (b) increasing $\mathbb{E}[R]$ does lead to better performance for the cycling estimator (but not for the simpler one), as expected. The CPU cost of cycling is less than 0.01 seconds higher than that of the simple estimator.
@@ 4.3.1. Variance Comparison (fixed $\theta$)
Table 2. Work-normalised variance and proportion of negative estimates out of $10^{3}$ replicates of $\widehat{\mathcal{Z}(\theta)^{-1}}$ for the simple and the cycling estimator when the variance of the estimates of $\mathcal{Z}(\theta)$ is moderate and low. Bold denotes the best values.
4.3.2. Posterior inference and model choice
To recover the posterior distribution and the marginal likelihood, we use importance sampling, with a Gaussian proposal $q$, from which we generate $n=1024$ vectors $\theta_{i}$ through randomised quasi-Monte Carlo. Using quasi-Monte Carlo ensures that the variability of our importance sampling estimates is dominated by the randomness of the estimates of $1/\mathcal{Z}(\theta)$. (See the appendix for more details on the proposal and on randomised quasi-Monte Carlo.)
@@ 4.3.2. Posterior inference and model choice
For each $\theta_{i}$, we use exactly the same approach (to estimate $1/\mathcal{Z}(\theta_{i})$) as the one which led to best performance in the previous section, namely, $n_{0}=10$, $\mathbb{E}[R]=10$, and waste-free SMC run in the “moderate variance” regime.
@@ 5. Conclusion
We develop a general strategy for unbiasedly estimating $f(m)$, for a smooth function $f$, from random variables $X_{i}$ with expectation $m$. Inspired by the simple estimator (Section 2.2) – which is widely used in the literature – we proposed a new cycling estimator (Section 2.3), and obtained conditions that guarantee finite estimation and execution time (CPU cost) variance (for both methods.) This is an improvement over previous approaches. In particular, the multi-level Monte Carlo method of Luo et al., (2020) returns estimates with infinite variance, while the one of Shi and Cornish, (2021) has a running time with infinite variance. The latter issue may be particularly problematic in a parallel implementation since the completion time is the maximum of all threads being run.
@@
We show that the cycling estimator outperforms the simple one both theoretically (Corollary 10 and Example 8) and empirically (Section 4), with potentially only a little more computational cost. We also discussed the sensitivity of these methods to their tuning parameters, an issue we found to be largely overlooked in the literature. We gave precise guidelines for selecting the expansion point $x_{0}$ and the distribution of the random variable $R$ that determines the random truncation (Sections 3.1 and 3.2).
@@
Our experiments show that, in order to build reliable unbiased estimates of $f(m)$ (or its gradient), one needs input variables $X_{i}$ with low or moderate variance. In fact, the minimal $\beta^{2}$ in Proposition 7, achieved for $x_{0}^{\star}$ in (16), is $(\beta^{2})^{\star}=\sigma^{2}/(\sigma^{2}+m^{2})$. (Note that $1-(\beta^{2})^{\star}$ is indeed the inverse of relative variance of the $X_{i}$’s.) As $\sigma^{2}/m^{2}\rightarrow\infty$, we need to select $p\in(0,1-(\beta^{2})^{\star})\to\{0\}$, hence requiring an infinite computational cost. In this case, we recommend dedicating the computational effort towards building lower variance estimates of $m$.
@@
Even if the $X_{i}$’s have low variance, a poorly chosen $x_{0}$ can cause the estimators of $f(m)$ to have high variance. While in low dimensional problems one can get away with a poorly tuned $x_{0}$, we found that for high dimensional applications (e.g. the ICA example in Section 4.2) it is crucial that the input variables $X_{i}$’s have low variance and to fine tune $x_{0}$, as recommended in Section 3.1. Once $x_{0}$ has been selected, we describe a simple strategy to select $p$ to guarantee that $\mathbb{E}[R]$ is sufficiently large to have finite and small variance. This strategy works well for the experiments we considered, however, in some cases it might be possible to build optimal distributions for particular $f$’s (see, e.g. Rhee and Glynn, (2015); Cui et al., (2021); Zheng et al., (2023)).
@@ Acknowledgments
Research of the first two authors is partly supported by the French Agence Nationale de la Recherche (ANR) through grant EPISVACAGE. SSS holds the Tibra Foundation professorial chair and is also supported by the Air Force Office of Scientific Research under award number FA2386-23-1-4100. We would like to thank Otmane Sakhi, Adrien Corenflos and Sam Power for helpful comments at various stages of this project.
@@ A.1. Proof of Lemma 2
(21) $\mathbb{E}\left[\sum_{k=0}^{\infty}\frac{\mathds{1}\{R\geq k\}}{\mathbb{P}(R\geq k)}|\gamma_{k}|U_{R,k}|\right]<\infty,$
@@
$\mathbb{E}\left[\sum_{k=0}^{\infty}\frac{\mathds{1}\{R\geq k\}}{\mathbb{P}(R\geq k)}|\gamma_{k}|U_{R,k}|\right]$ $=\sum_{k=0}^{\infty}\frac{\mathbb{E}\left[\mathds{1}\{R\geq k\}|U_{R,k}|\right]}{\mathbb{P}(R\geq k)}|\gamma_{k}|\leq\sum_{k=0}^{\infty}|\gamma_{k}|a_{k}<\infty,$
@@ A.3. Verification of (7)
$\mathbb{E}[\hat{f}|R]^{2}=$ $\sum_{k=0}^{\infty}\gamma_{k}^{2}\frac{\mathds{1}\{R\geq k\}}{\mathbb{P}(R\geq k)^{2}}\left(\frac{m}{x_{0}}-1\right)^{2k}$
$+2\sum_{k=0}^{\infty}\sum_{l=k+1}^{\infty}\gamma_{k}\gamma_{l}\frac{\mathds{1}\{R\geq l\}}{\mathbb{P}(R\geq k)\mathbb{P}(R\geq l)}\left(\frac{m}{x_{0}}-1\right)^{k+l}.$

### v3

@@ 4.1. Toy latent variable model
We set $\theta=1$ and sample a data set of size $n=10^{3}$. The goal is to unbiasedly estimate the log-likelihood $\sum_{i=1}^{n}\log p(y_{i}|\theta)$ for $\theta=1$, a quantity which we are able to compute exactly for this model.
@@
$I_{j}:=\log\frac{1}{2^{j}}\sum_{h=1}^{2^{j}}w(Z_{h};y,\theta),\quad\text{with }w(Z_{h};y,\theta):=\frac{p(Z_{h},y|\theta)}{q(Z_{h};y)},\quad Z_{1},\dots,Z_{2^{j}}\sim q(\cdot;y).$
and $\Delta_{k}:=I_{j+k+1}-(I_{j+k}^{E}+I_{j+k}^{O})/2$; with $I_{j+k}^{E}$ (resp. $I_{j+k}^{O}$) being the IWAE estimate that only uses the even-indexed (resp. odd-indexed) samples. We consider the same proposal as in Shi and Cornish, (2021), i.e. $\mathcal{N}(\cdot;(y+\theta^{\star}\textsf{1}_{d})/2,\frac{2}{3}\textsf{Id}_{d})$, where $\theta^{\star}$ denotes the maximum likelihood estimator for $\theta$, which is shown to be optimal within a certain family in Rainforth et al., (2018).
The expected CPU cost, i.e. the number of generated $X_{j}$, is $\mathbb{E}[R]=1/p-1$ for our Taylor-based estimators, and $\mathbb{E}[2^{j+1+\widetilde{R}}]$ for the MLMC approach of Shi and Cornish, (2021). We adjust $p$ and $j$ so that both expectations are equal to a certain pre-specified budget $C$. We consider two values for $C$, namely $C=6$ and $C=96$.
@@
This experiment shows that the cost of MLMC has large empirical variance. In fact, the true variance of the cost is infinite, see Appendix E.1 for a proof of this statement. In contrast, the cost of the simple and cycling estimators remains close to its expectation.
The simple and cycling estimators outperform MLMC in terms of accuracy, with the simple estimator having larger variance than the cycling estimator for higher computational cost. This is particularly evident in the first plot of the second row, in which we compare the estimates for $\mathbb{E}[R]=96$, in this case the cloud of estimates obtained with cycling is considerably more concentrated around the true value (the horizontal line). This is consistent with the theoretical results established in Section 2 which show that the variance of the simple estimator does not decay to 0 when the computational cost increases.
@@
As a note of caution, our Taylor-based estimators may give very poor results for a larger dimension $d$; see Appendix E.1 for extra results when $d=20$. Simple calculations show that, in that case, one would need to take $\mathbb{E}[R]\gg 10^{8}$ to ensure that $p<1-\beta^{2}$ and that the estimators have finite variance. In this case, the inputs have very large variance, and presumably the only way to fix this issue is to come up with a different strategy to unbiasedly estimate $p(y|\theta)$, so that the corresponding estimates have much lower variance.
@@ 4.2. Independent component analysis
To estimate $\theta$, Allassonnière and Younes, (2012) use a stochastic approximation expectation maximisation (SAEM) algorithm. We will instead use stochastic gradient descent (Robbins and Monro,, 1951), implemented with our unbiased estimators of $\nabla_{\theta}f(m(\theta))=\nabla_{\theta}\log p(\theta|y)$. We call our two implementations simple-SGD and cycling-SGD. We also implement SGD with the gradient unbiasedly estimated using the MLMC strategy of Shi and Cornish, (2021) (MLMC-SGD). The bivariates $(X_{i},G_{i})$ (for $i=1,2,\ldots$) of Section 2.6 are generated using importance sampling; and the same importance sampling proposal is used for MLMC. For MLMC, we adjust its parameters $p$ and $j$, defined in (18) with $\tilde{R}\sim\mathrm{Geometric}(p)$, so that $\mathbb{E}\left[2^{j+1+\widetilde{R}}\right]=n_{0}$. Similarly, $p$ for the Geometric random truncation of the simple and cycling gradient estimates is found by setting $\mathbb{E}\left[R\right]=n_{0}$. See Appendix E.3 for all the details.
@@
SAEM $7.00\cdot 10^{-7}$ 0.22 33
MLMC $2.95\cdot 10^{-6}$ 0.29 12 19311 11
simple $2.97\cdot 10^{-6}$ 0.27 10 90 10
cycling $6.80\cdot 10^{-5}$ 0.24 10 90 14
@@
We also observe that there seems to be a small difference between the results obtained with the simple and the cycling estimator. Both are unbiased, however, while running our experiments we noticed that the higher variance of the simple estimator requires smaller learning rates to counterbalance the gradient estimate.
@@ 4.3. Exponential random graphs
We construct an unbiased estimate of $1/\mathcal{Z}(\theta)$, for any $\theta$, which may then be combined with the remaining tractable terms of (20) to perform Bayesian inference and model choice. We consider the same example as in Caimo and Friel, (2011): an $n=16$ node graph; the statistic is $s(y)=(\sum_{i<j}y_{ij},\sum_{i<j<k}y_{ij}y_{jk})$, which counts the number of edges and of two stars; $\theta\in{}^{2}$ is assigned the prior $N_{2}(0_{2},30I_{2})$. The data – the observed connectivity – is the Florentine family business graph (Padgett and Ansell,, 1993).
@@
We first fix $\theta=(-3.15,0.58)$, a value in the high probability region of the posterior, and compare the performance of the simple and the cycling estimator.
@@
Table 2. Work-normalised variance and proportion of negative estimates out of $10^{3}$ replicates of $\widehat{\mathcal{Z}(\theta)^{-1}}$ for the simple and the cycling estimator when the variance of the estimates of $\mathcal{Z}(\theta)$ is moderate and low.
We select the value of $x_{0}$ using the strategy in Section 3.1 with $n_{0}=10$. Once $x_{0}$ has been selected, we identify the corresponding $\widehat{\beta^{2}}$ as discussed in Section 3.2, see Algorithm 1 for a summary. This gives us a range for $p\in(0,1-\widehat{\beta^{2}})$. In the moderate variance regime we obtain $p\in(0,0.7)$ while for the low variance regime we have $p\in(0,0.95)$. In both cases, it is sufficient to set $\mathbb{E}[R]=1$ (i.e. $p=0.5$) to ensure that both the simple and the cycling estimator have finite variance.
Since the cost of obtaining the estimates of $\mathcal{Z}(\theta)$ is significantly larger than that of computing the estimates, selecting $\mathbb{E}[R]=1$ might be wasteful as our theoretical results in Section 2 guarantee that the cycling estimator has a lower variance than the simple estimator for large values of $\mathbb{E}[R]$. Thus, we also consider $\mathbb{E}[R]=10=n_{0}$ in our simulations as recommended in Section 3.2.
Table 2 compares the simple and cycling estimators in these four different scenarios (two variance regimes, two values of $\mathbb{E}[R]$) according to two metrics estimated on $10^{3}$ independent runs: work-normalised variance (relative variance times CPU cost), and proportion of negative $1/\mathcal{Z}(\theta)$ estimates. The results show that (a) the cycling estimator systematically outperforms the simple one; and (b) increasing $\mathbb{E}[R]$ does lead to better performance for the cycling estimator but not for the simpler one, as expected. The CPU cost of cycling is less than 0.01 seconds higher than that of the simple estimator.
To recover the posterior distribution and the marginal likelihood, we use importance sampling, with a Gaussian proposal $q$, from which we generate $n=1024$ vectors $\theta_{i}$ through randomised quasi-Monte Carlo. Using quasi-Monte Carlo ensures that the variability of our importance sampling estimates is dominated by the randomness of the estimates of $1/\mathcal{Z}(\theta)$. See the appendix for more details on the proposal and on randomised quasi-Monte Carlo.
@@
For each $\theta_{i}$, we use exactly the same approach to estimate $1/\mathcal{Z}(\theta_{i})$ as the one which led to best performance in the previous section, namely, $n_{0}=10$, $\mathbb{E}[R]=10$, and waste-free SMC run in the “moderate variance” regime.
@@ 5. Conclusion
We develop a general strategy for unbiasedly estimating $f(m)$, for a smooth function $f$, from random variables $X_{i}$ with expectation $m$. Inspired by the simple estimator (Section 2.2) – which is widely used in the literature – we proposed a new cycling estimator (Section 2.3), and obtained conditions that guarantee finite estimator and execution time (CPU cost) variance for both methods. This is an improvement over previous approaches. In particular, the multi-level Monte Carlo method of Luo et al., (2020) returns estimates with infinite variance, while the one of Shi and Cornish, (2021) has a running time with infinite variance. The latter issue may be particularly problematic in a parallel implementation since the completion time is the maximum of all threads being run.
As pointed out by a referee, the fact that the execution cost is random may be a drawback in certain implementation scenarios, even if it has finite variance. In such a case, one may prefer using a biased estimate whose execution cost is deterministic. Such an approach has less theoretical support, but seems to work well in certain practical applications, e.g., to approximate gradients in variational auto-encoders (Burda et al.,, 2016).
@@
We show that the cycling estimator outperforms the simple one both theoretically, see Proposition 9 and Example 8 and empirically, see Section 4, with potentially only a little more computational cost. We also discussed the sensitivity of these methods to their tuning parameters, an issue we found to be largely overlooked in the literature. We gave precise guidelines for selecting the expansion point $x_{0}$ and the distribution of the random variable $R$ that determines the random truncation (Sections 3.1 and 3.2).
@@
Our experiments show that, in order to build reliable unbiased estimates of $f(m)$ or its gradient, one needs input variables $X_{i}$ with low or moderate variance. In fact, the minimal $\beta^{2}$ in Proposition 7, achieved for $x_{0}^{\star}$ in (17), is $(\beta^{2})^{\star}=\sigma^{2}/(\sigma^{2}+m^{2})$. Note that $1-(\beta^{2})^{\star}$ is indeed the inverse of relative variance of the $X_{i}$’s. As $\sigma^{2}/m^{2}\rightarrow\infty$, we need to select $p\in(0,1-(\beta^{2})^{\star})\to\{0\}$, hence requiring an infinite computational cost. In this case, we recommend dedicating the computational effort towards building lower variance estimates of $m$.
@@
Even if the $X_{i}$’s have low variance, a poorly chosen $x_{0}$ can cause the estimators of $f(m)$ to have high variance. While in low dimensional problems one can get away with a poorly tuned $x_{0}$, we found that for high dimensional applications, e.g., the ICA example in Section 4.2, it is crucial that the input variables $X_{i}$’s have low variance and to fine tune $x_{0}$, as recommended in Section 3.1. Once $x_{0}$ has been selected, we describe a simple strategy to select $p$ to guarantee that $\mathbb{E}[R]$ is sufficiently large to have finite and small variance. This strategy works well for the experiments we considered, however, in some cases it might be possible to build optimal distributions for particular $f$’s (see, e.g. Rhee and Glynn, (2015); Cui et al., (2021); Zheng et al., (2023)).
As a by-product of our study, we provide a method to unbiasedly estimate the model evidence in doubly intractable models. Previous attempts in the context of ERG models have focused on pseudo-likelihood approaches in which the intractable likelihood is replaced by a tractable approximation (Bouranis et al.,, 2018). Also, using a combination of path sampling and kernel density estimation, which provides biased estimators and are only suitable in low-dimensional settings (Caimo and Friel,, 2013, Section 4).
@@ Acknowledgments
Research of the first two authors is partly supported by the French Agence Nationale de la Recherche (ANR) through grant EPISVACAGE. SSS holds the Tibra Foundation professorial chair and gratefully acknowledges research funding as follows: This material is based upon work supported by the Air Force Office of Scientific Research under award number FA2386-23-1-4100. The authors would like to thank Otmane Sakhi, Adrien Corenflos Sam Power and two referees for helpful comments at various stages of this project.
@@ A.1. Proof of Lemma 2
(22) $\mathbb{E}\left[\sum_{k=0}^{\infty}\frac{\mathds{1}\{R\geq k\}}{\mathbb{P}(R\geq k)}|\gamma_{k}||U_{R,k}|\right]<\infty,$
@@
$\mathbb{E}\left[\sum_{k=0}^{\infty}\frac{\mathds{1}\{R\geq k\}}{\mathbb{P}(R\geq k)}|\gamma_{k}||U_{R,k}|\right]$ $=\sum_{k=0}^{\infty}\frac{\mathbb{E}\left[\mathds{1}\{R\geq k\}|U_{R,k}|\right]}{\mathbb{P}(R\geq k)}|\gamma_{k}|\leq\sum_{k=0}^{\infty}|\gamma_{k}|a_{k}<\infty,$
@@ A.3. Verification of (7)
(25) $\mathbb{E}[\hat{f}|R]^{2}=$ $\sum_{k=0}^{\infty}\gamma_{k}^{2}\frac{\mathds{1}\{R\geq k\}}{\mathbb{P}(R\geq k)^{2}}\left(\frac{m}{x_{0}}-1\right)^{2k}$
(26) $+2\sum_{k=0}^{\infty}\sum_{l=k+1}^{\infty}\gamma_{k}\gamma_{l}\frac{\mathds{1}\{R\geq l\}}{\mathbb{P}(R\geq k)\mathbb{P}(R\geq l)}\left(\frac{m}{x_{0}}-1\right)^{k+l}.$
@@
A.4. Derivation of (9)
We start by computing $\mathbb{E}[(\hat{f})^{2}|R=r]$:
$\mathbb{E}[(\hat{f})^{2}|R=r]$ $=\mathbb{E}\left[\left(\sum_{k=0}^{\infty}\frac{\mathds{1}\{R\geq k\}}{\mathbb{P}(R\geq k)}\gamma_{k}U_{R,k}\right)^{2}|R=r\right]$
$=\mathbb{E}\left[\sum_{k=0}^{\infty}\frac{\mathds{1}\{R\geq k\}}{\mathbb{P}(R\geq k)^{2}}\gamma_{k}^{2}U_{R,k}^{2}|R=r\right]$
$+\mathbb{E}\left[2\sum_{k=0}^{\infty}\sum_{l=k+1}^{\infty}\frac{\mathds{1}\{R\geq k\}\mathds{1}\{R\geq l\}}{\mathbb{P}(R\geq k)\mathbb{P}(R\geq l)}\gamma_{k}\gamma_{l}U_{R,k}U_{R,l}|R=r\right]$
$=\sum_{k=0}^{\infty}\frac{\mathds{1}\{r\geq k\}}{\mathbb{P}(R\geq k)^{2}}\gamma_{k}^{2}U_{r,k}^{2}+2\sum_{k=0}^{\infty}\sum_{l=k+1}^{\infty}\frac{\mathds{1}\{r\geq l\}}{\mathbb{P}(R\geq k)\mathbb{P}(R\geq l)}\gamma_{k}\gamma_{l}U_{r,k}U_{r,l}.$
This and (25) give
$\operatorname{var}[\hat{f}|R=r]$ $=\sum_{k=0}^{\infty}\frac{\mathds{1}\{r\geq k\}}{\mathbb{P}(R\geq k)^{2}}\gamma_{k}^{2}\left(U_{r,k}^{2}-\left(\frac{m}{x_{0}}-1\right)^{2k}\right)$
$+2\sum_{k=0}^{\infty}\sum_{l=k+1}^{\infty}\frac{\mathds{1}\{r\geq l\}}{\mathbb{P}(R\geq k)\mathbb{P}(R\geq l)}\gamma_{k}\gamma_{l}\left(U_{r,k}U_{r,l}-\left(\frac{m}{x_{0}}-1\right)^{k+l}\right)$
$=\sum_{k=0}^{\infty}\frac{\mathds{1}\{r\geq k\}}{\mathbb{P}(R\geq k)^{2}}\gamma_{k}^{2}\operatorname{var}(U_{r,k})$
$+2\sum_{k=0}^{\infty}\sum_{l=k+1}^{\infty}\frac{\mathds{1}\{r\geq l\}}{\mathbb{P}(R\geq k)\mathbb{P}(R\geq l)}\gamma_{k}\gamma_{l}\operatorname{cov}\left(U_{r,k},U_{r,l}\right).$
Taking expectations w.r.t. $R$ gives the result.